# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v1

@@ 4.2 Experimental Results
Table 2: Main results of the models trained on Qwen2.5.(See Appendix J for FLOPs analysis)
@@
BERT-Based System 81.43% 38.79
Qwen2.5-72B (ICL) 94.74% 5223.91
Qwen2.5-7B (RFT) 95.29% 572.47
Qwen2.5-3B (RFT) 95.30% 297.98
Qwen2.5-1.5B (RFT) 95.16% 171.19
Qwen2.5-0.5B (RFT) 94.66% 77.89
Qwen2.5-0.5B (KD) 94.85% 77.89
@@
Qwen2.5-0.4B-Width (KD) 94.17% 72.93
Qwen2.5-0.4B-Depth (KD) 94.20% 67.68
@@
Qwen2.5-0.5B (RFT) 94.66% 89.32%
Qwen2.5-3B (RFT) 95.30% 91.02%
Qwen2.5-3B (RL) 94.68% 92.95%
@@
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over the crowdsource labeler. This demonstrates the superior performance of the LLM prototype. Using the LLM-as-a-Judge for evaluation, we find that the LLM Prototype surpasses the Bert-Based System by 13.31% and crowdsourced annotations by 16.89%, as displayed in Table 2. This conclusion is roughly consistent with those provided by professional annotators, demonstrating the effectiveness of the LLM judge.
@@
As demonstrated in Table 2, models trained with RFT generally exhibit a higher AR than the LLM prototype. Specifically, both the 3B and 7B models display negligible differences in performance but each achieve a 0.5% improvement in AR compared to the LLM prototype. In contrast, the 0.5B model shows a 0.5% lower AR than the 1.5B model. Moreover, KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%. In terms of inference performance, the 0.5B model demonstrates only a slight 0.4% difference in AR compared to the 7B model, while achieving a significant 13.57x improvement in QPS and a substantial 91.3% reduction in latency, as illustrated in Table 4.
@@
According to Table 3, while RFT performs well with in-domain data, it demonstrates poorer results with out-of-domain data, showing a 3.61% AR decrease compared to the LLM prototype. Conversely, RL models present a smaller gap (1.68%) with the LLM prototype on out-of-domain data, indicating better generalization of RL training methodologies.
Tiny LMs as a Cost-Efficient Upgrade to BERT-based System
Table 2 and Table 4 illustrate that quantization introduces a negligible AR loss of less than 0.1% while significantly enhancing inference performance. In the optimal scenario, quantization results in a 1.29x increase in QPS and a 0.76x decrease in RT. On the other hand, pruning results in a slightly higher accuracy loss, approximately 0.2%, but offers a more consistent 20% improvement in performance. In terms of end-to-end inference performance, under the condition where the CC is 64, the BERT-based system achieves a QPS of 1110.88 and an RT of 55.84 ms. Despite the cost being approximately doubled compared to the BERT-based system, the AR is improved by an absolute increase of 13%.
@@ 5 Analysis
Our experimental analysis highlights a compelling trade-off between model size and computational efficiency. Although larger models typically demonstrate superior performance, the expected disparity in performance between large and smaller models is not as pronounced as anticipated. As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs, while even reaches more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC. Although this model has a slight gap in cost compared to traditional BERT models, the attractive accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
Characteristics of Training Paradigm
The results presented in Table 3 underscore the distinct characteristics of different training methodologies. Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set distribution. In contrast, RL demonstrates greater robustness by not merely fitting individual data points. We posit that RL’s superior generalization capabilities stem from its ability to learn from negative samples, which are generated by the model being trained itself. RL is highly effective for dealing with well-structured problems, while it presents unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks. Moreover, RL is sensitive to noisy data. Despite these limitations, RL remains a promising direction for development, especially when building upon the foundation of traditional fine-tuning training strategies.
@@
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at equivalent compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning demonstrates optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms under high concurrency conditions with lossless accuracy, considered a preferred deployment paradigm, highlighting the importance of throughput.
@@ 6 Conclusion
Cost-Efficient Three-Stage Pipeline
The proposed pipeline—comprising prototyping, knowledge transfer, and model compression—presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference. We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
@@ 7 Limitation & Future Works
While our proposed three-stage pipeline demonstrates promising results in cost-performance, several limitations warrant further investigation, underscoring the need for holistic optimization across algorithmic robustness, dynamic adaptation and hardware efficiency to enable truly scalable and trustworthy adoption in industrial settings:
Continual Learning for Dynamic Adaptation
While cost-efficient, the static training-deployment cycle of our pipeline hinders its adaptation to dynamic tasks and evolving datasets. Future work should develop advanced continual learning frameworks that enable incremental knowledge integration without catastrophic forgetting, leveraging rapid training cycles to maintain deployment agility. Employing online learning and lifelong learning will enhance models’ robustness and adaptability, ensuring sustained relevance and precision in changing linguistic contexts and shifting user demands.
Edge Deployment Optimization with Performance Assurance
Being sufficiently compact, our super tiny models have demonstrated impressive capabilities in currently high-demand applications of IoT devices, such as mobile phones and tablets, characterized by low latency, cost-effectiveness, cloud independence, and robust privacy. While further compression to an even smaller size like 0.1B may not be significantly beneficial, exploration of optimizing on-device deployment of such ultra-small models remains a promising direction for future research. Future works should develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, addressing real-world heterogeneous hardware constraints such as memory fragmentation and NPU compatibility. We will also explore more resource-efficient deployment techniques for edge devices—such as sparse computation, edge-centric quantization, and federated learning.
@@ Appendix A Dataset Analysis
Type-Token Rate (TTR) is a fundamental measure of lexical diversity that reflects the variability of vocabulary locally, but it is sensitive to text length. The calculation formula is:
@@
Entropy is an information-theoretic metric that quantifies the uncertainty within binary grammatical sequences, based on their probability distribution. A higher entropy value indicates more complex local language patterns. The calculation formula is:
@@
Both TTR and Entropy are sensitive to text length. To enable a relatively fair comparison of the complexity across datasets, we randomly sample 700 entries from the test set of each dataset for the computation(as the ATIS dataset only contains 700 instances).
@@ Appendix B Details for Reinforcement Learning
  • For Format Reward, we adopt the logic of Logic-RL[9], which utilizes regular expressions to validate the output format against the specified template <think>...</think><answer>...</answer>. Additionally, we extend this design by verifying the legality of function usage and its parameters. For example, if it choose a function that doesn’t exist in the input function pool, then it is considered as invalid.
@@
  • For Answer Reward, according to our specific scenario, we find that the model attempts to directly guess answer during the training, leading no association between the Chain-of-Thought (CoT) and the final results. Requiring explicit repetition of the answer within the CoT framework effectively resolves this issue. As demonstrated in Figure 6, we optimize its original design in order to make the Chain-of-Thought (CoT) reasoning and the results more consistent. We require: (1) the CoT must explicitly reference the target function at least once; (2) the final function mentioned in the CoT must match the function in the answer. Otherwise, it will be considered as a parsing error.
$S_{answer}=\begin{cases}2,&\text{if the answer is correct,}\\ &\text{and the thought ends with the function in the answer,}\\ -2&\text{if the answer cannot be parsed or is missing,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.5,&\text{otherwise}\end{cases}$ (10)
@@ Appendix D Retriever Setup
Initially, we employ GTE-large [59] as the embedding model trained based on the BERT-large [71]. We compare the cosine similarity of the embedding between the API description generated by LLM based on this model and the API description in the API document, but the recall accuracy is not high.
Subsequently, we replace it with GTE-Qwen2-1.5B-instruct [59] trained based on Qwen2-1.5B [97]. This model has the advantage of being able to fine tune matching tasks by modifying prompts.
Next, based on this model and a customized prompt in Section D.1, we achieve a higher accuracy by matching the cosine similarity between the API description generated by the model and the JSON text of the entire API document. During training, we optimize the API description model through rejection sampling, with implementation details provided in Section D.2.
Finally, to achieve maximal inference speed, we trained a classifier based on BERT-large [71] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and outputs the function category. Then, the functions corresponding to the top-k output logits are considered as candidate functions for the subsequent model’s function call. This approach significantly improves inference speed (due to the inherent efficiency of BERT classifiers compared to autoregressive GPT models) and also enhances function recall. However, this comes at the cost of scalability, as retraining the recall model is required for each update to the function pool.
As illustrated in Table 5, these optimizations substantially improve API recall rate.
@@ D.1 Prompt for retrieval
Instruction: Given a API description, retrieve the most similar API from the function pool
@@ Appendix E Training Details
Table 6: Training Parameters in different stages.
@@
The main training parameters are shown in Table 7. During the training phase, we utilize the AdamW optimizer [96], setting the model’s maximum sequence length to $5120$. In the RL phase, we employ the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages of efficiency, conciseness, and training stability. Following [121], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. For pruning, we use the 0.5B-Instruct model as the base model and employ width/depth pruning strategies with 20%/30% prune ratio, finally obtaining two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. We produce the extremely compressed models through a two-stage training process (accuracy recovery training, KD). Complete training details for the pruned models can be found in Appendix I. For quantization, we primarily experiment with three quantization strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 adopt per-group symmetric int8 weight-only quantization with a group_size=128, and activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 applies FP8 for both weights and activations. For the training frameworks, we use Megatron-LM [73] for SFT and KD, while OpenRLHF [74] is employed for RL. Data sampling during training is implemented using vLLM [72]. In the evaluation phase, we utilize SGLang [75] to generate data.
@@ Appendix F Computing resources
We utilize 8 NVIDIA H20 GPUs for training process and data generation. For inference performance evaluation, we employ a single NVIDIA L20 GPU. We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, which operates at a frequency of 3028.956 MHz and a cache size of 327680 KB. In addition, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory.
During the training phase, the RFT process takes approximately 2 to 14 hours for models ranging from 0.5B to 7B. Pre-KD logits caching for a 7B model requires about 5 hours, while RL for the 3B model takes approximately 3 days. The KD process for the 0.5B model is completed in about 2 hours. Training each pruned model demands roughly 3.5 days.
Furthermore, for the LLM prototype and LLM-as-a-Judge, data generation of every 10,000 entries requires approximately 1 hour.

### v2

@@ 4.2 Experimental Results
Table 2: Main results of the models trained on Qwen2.5(See Appendix J for FLOPs analysis).
@@
BERT-based System 81.43% 38.65
Qwen2.5-72B (ICL) 94.74% 4907.50
Qwen2.5-7B (RFT) 95.29% 434.04
Qwen2.5-3B (RFT) 95.30% 218.94
Qwen2.5-1.5B (RFT) 95.16% 111.91
Qwen2.5-0.5B (RFT) 94.66% 43.31
Qwen2.5-0.5B (KD) 94.85% 43.31
@@
Qwen2.5-0.4B-Width (KD) 94.17% 33.10
Qwen2.5-0.4B-Depth (KD) 94.20% 40.83
@@
Qwen2.5-7B (RFT) 95.29% 90.86%
Qwen2.5-0.5B (RFT) 94.66% 90.62%
Qwen2.5-7B (RL) 95.36% 94.80%
Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
@@
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a win+tie rate of 95.59% over the BERT-based system and 92.98% over crowdsourced labelers. This demonstrates the superior performance of the LLM prototype. Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based System by 13.31% and crowdsourced annotations by 16.89%, as shown in Table 2. This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
@@
As demonstrated in Table 2, models trained with RFT generally exhibit a higher AR than the LLM prototype. Specifically, both the 3B and 7B models display negligible differences in performance and both achieves a 0.5% improvement in AR compared to the LLM prototype. The 0.5B model’s AR is 0.5% lower than the 1.5B model, while KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%. In terms of inference performance, the 0.5B model demonstrates only a slight 0.4% difference in AR compared to the 7B model, while achieving a significant 13.57x improvement in QPS and a substantial 91.3% reduction in latency, as illustrated in Table 4.
@@
According to Table 3, while RFT performs well on in-domain data, it demonstrates worse results on out-of-domain data with a 3.61% AR decrease compared to the LLM prototype. Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology. Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1% in out-of-domain datasets compared to the RFT method. For more detailed exploration of the RL results, please refer to Appendix K.
Tiny LLMs as a Cost-Efficient Upgrade to BERT-based System
As shown in Table 2 and Table 4, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario. Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%. For end-to-end inference, with a CC of 64, the BERT-based system reaches 1110.88 QPS and 55.84 ms RT. Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
@@ 5 Analysis
Our experimental analysis highlights a compelling trade-off between model size and computational efficiency. Although larger models typically show superior performance, the expected disparity between large and small ones is not as pronounced as anticipated. As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs and reaching more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC. Although this model has a slight gap in cost compared to traditional BERT models, the improved accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
Characteristics of Training Paradigms
The results presented in Table 3 underscore the distinct characteristics of different training methodologies. Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set. In contrast, RL presents greater robustness by not merely fitting individual data points. We posit that RL’s superior generalization capabilities stem from its ability to learn from negative samples, which are generated by the model being trained itself. RL is highly effective for dealing with well-structured problems, while it has unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks. Additionally, applying RL directly to tiny models often proves to be ineffective, likely due to such models’ limited capacity to utilize Chain-of-Thought (CoT) rationales. In such cases, KD can compensate for this limitation by training the tiny models on a large amount of CoT data generated by a RL-trained teacher model. Compared to fine-tuning, KD can better extract CoT rationales from the probability distribution. Consequently, the hybrid methodology of KD and RL can achieve remarkably enhanced performance and represents a promising direction for future development.
@@
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 2), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 9), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix J). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
@@ 6 Conclusion
Scalable Cost-Efficient Three-Stage Pipeline
The proposed pipeline-comprising prototyping, knowledge transfer, and model compression-presents a systematic framework designed to achieve high performance while maintaining low costs in both training and inference. We believe that this approach can be effectively adapted to other domains, thereby potentially advancing research in those areas and facilitating the widespread deployment of LLMs.
@@ 7 Limitation & Future Works
Our pipeline has presented optimal cost-performance results, and our super tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy. Further compression may offer limited marginal benefits. Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications such as mobile devices:
Hybrid Methodology for LLMs Capabilities Transfer
Based on our existing experiments, the mixer of reinforcement learning and knowledge distillation has effectively enhanced the generalization capabilities of super tiny models. This implies a potentially powerful hybrid methodology that allows super tiny models to nearly fully retain the capabilities of LLMs in domain-specific tasks. In future research, we will build upon our previous work to conduct deeper studies and theoretical analyses of this hybrid approach in transferring LLMs capabilities to such super tiny models.
On-device Deployment Optimization
Although we achieve 8-bit compression and online deployment, real-world edge devices face heterogeneous hardware constraints such as memory fragmentation and NPU compatibility. In the future, for on-device deployment of super tiny models, we will develop hardware-aware compression algorithms and adaptive runtime kernels to optimize latency-energy tradeoffs across diverse IoT ecosystems, and explore more resource-efficient deployment techniques for edge devices, such as sparse computation, edge-centric quantization, and federated learning.
@@ Appendix A Dataset Analysis
Type-Token Rate (TTR), as a fundamental measure of lexical diversity, reflects the local variability of vocabulary, but it is sensitive to text length. The calculation formula is:
@@
Entropy, as an information-theoretic metric, quantifies the uncertainty within binary grammatical sequences based on their probability distribution. A higher entropy value indicates more complex local language patterns. The calculation formula is:
@@
Since both TTR and Entropy are sensitive to text length, to ensure a relatively fair comparison of dataset complexity, we randomly sample 700 entries from the test set of each dataset for calculation(as the ATIS dataset only contains 700 instances).
@@ Appendix B Details for Reinforcement Learning
  • For Format Reward, we adopt the logic of Logic-RL[9], which validates the output format against the specified template <think>...</think><answer>...</answer> via regular expressions. We add additional designs to ensure the validity of the usage of the function and its parameters. If a function not present in the predefined input function pool is selected, it is deemed invalid.
@@
  • For Answer Reward, in our specific scenario, we observe that the model tends to directly guess answer during training, resulting in no association between the Chain-of-Thought (CoT) and the final results. Requiring explicit repetition of the answer within the CoT framework effectively resolves this issue. As illustrated in Figure 6, we optimize its original design in order to enhance the consistency between the Chain-of-Thought (CoT) reasoning and the results. Our requirements are as follows: (1) The CoT must explicitly reference the target function at least once; (2) The final function mentioned in the CoT must correspond to the function in the answer. Failure to comply with these conditions will be regarded as a parsing error. Additionally, we incorporate fine-grained rewards, which provide better guidance when the answer is partially correct, enabling the model to achieve improved performance.
$S_{answer}=\begin{cases}-2&\text{if the answer is missing, or is unable to be parsed,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.25&\text{if the names of at least one parameter and function are correct,}\\ -1.2&\text{if the name of function, as well as the name }\\ &\text{and value of at least one parameter are correct}\\ 2,&\text{if the answer is correct,}\\ -1.5,&\text{otherwise}\end{cases}$ (10)
@@ Appendix D Retriever Setup
Initially, we select GTE-large [63], which is trained on the BERT-large [75], as our embedding model. Using this model, we generate embeddings of API descriptions produced by the LLMs. Then, we compare these embeddings with those of the API description from the document using cosine similarity. However, this model’s recall accuracy is not high. Subsequently, we switch to GTE-Qwen2-1.5B-instruct [63], a model trained on Qwen2-1.5B [76], which offers the advantage of fine-tuning for matching tasks through prompt modification. Next, utilizing this model with a customized prompt in Section D.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy. During training, we optimize the API description model using rejection sampling, with implementation details provided in Section D.2. Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [75] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and the function category as output. The functions corresponding to the top-k output logits are then considered as candidate functions for the subsequent model’s function call. This approach significantly accelerates inference due to the efficiency of BERT classifiers compared to autoregressive GPT models, while also enhancing function recall capabilities. However, it introduces scalability challenges since the recall model needs to be retrained with each update to the function pool. As illustrated in Table 5, this series of optimizations lead to a substantially improvement in API recall rates.
@@ D.1 Prompt for retrieval
Instruction: Given a API description, retrieve the most similar API from the function pool.
@@ Appendix E Training Details
Table 6: Training Parameters in different phases.
@@
The main training parameters are detailed in Table 7. During the training phase, we employ the AdamW optimizer [77], setting the model’s maximum sequence length to $5120$. In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability. Following [78], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD). Comprehensive training details for the pruned models are available in Appendix I. For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 employ per-group symmetric int8 weight-only quantization with a group_size=128, while activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 uses FP8 for both weights and activations. For the training frameworks, we utilize Megatron-LM [79] for SFT and KD, while OpenRLHF [80] is employed for RL. Data sampling during training is conducted using vLLM [81]. In the evaluation phase, we employ SGLang [82] to generate data.
@@ Appendix F Computing resources
We utilize 8 NVIDIA H20 GPUs for training and data generation processes. For evaluating inference performance, a single NVIDIA L20 GPU is employed. We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, operating at a frequency of 3028.956 MHz and a cache size of 327680 KB. Additionally, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory. During the training phase, the RFT process spans approximately 2 to 14 hours for models ranging from 0.5B to 7B in size. Pre-KD logits caching for a 7B model requires about 5 hours, whereas RL for the 3B model extends to approximately 3 days. The KD process for the 0.5B model is completed within about 2 hours, and training each pruned model requires roughly 3.5 days. Moreover, for the LLM prototype and LLM-as-a-Judge, generating data for every 10,000 entries takes approximately 1 hour.